# Towards Causal Market Simulators

### v3

@@ 1. Introduction
Market generators (kondratyev2019market) are numerical techniques that rely on generative models for the purpose of synthetic market data generation. They leverage architectures such as generative adversarial networks (GANs) (wiese2020quant; li2020generating). For example, in market making, probabilistic forecasts can enhance trading strategies, allowing more profitable and risk-aware trading (lalor2025event). Important attributes of synthetic financial time series are the preservation of stylized facts (cont2001empirical) and rough path signatures (muca2024theoretical).
Historically, understanding of causality in times series was limited to Granger causality (kleinberg2009the). Granger causality assumes that the frequency of data measurement matches the true causal frequency of the underlying physical process (gong2017causal). However, making investment decisions requires a causal attribution of the premia to risk factors (lopezdeprado2025ai). As such, research begins to incorporate causal reasoning and explainability into time series foundation models (e.g., through attention mechanisms) (marconi2025tsfm; robertson2025dopfn). For example, (sokolov2025toward) demonstrated the utility of specifying and analyzing detailed causal models for financial markets for investment management. (oliveira2024causality) investigated superior financial time series forecasting that leverages causality-inspired models to balance the trade-off between invariance to distributional changes and minimization of prediction errors.
Research in generative models went beyond vector auto regression (VAR) (stock2001vector) to account for structural causality through the framework of structural causal models (SCMs) (pearl2009causality). Examples of time series causal methods are SCIGAN (bica2020estimatingthe), Time Series Deconfounder (bica2020time), and Causal Transformer (melnychuk2022causal). Deep structural causal models (DSCMs) (poinsot2024learning) enable a specific type of conditioned generation, counterfactuals, which are possible realistic scenarios that have not yet happened (horvath_generative_2025). They are a promising method for promoting financial stress tests (gao2018causal), risk management, scenario analysis, and backtesting (harvey2015backtesting).
@@ 2. Methodology
We propose a causal market simulator building on top of previous research in neural causal models (NCMs) (xia2023neural) and variational autoencoder (VAE) (kingma2013auto) based causal representation learning (scholkopf2021toward) in time series. Our neural causal model for time series (TNCM-VAE) [^1] https://github.com/thummd/tncm consists of three main components: an encoder to infer latent representations, a causal mapping module to handle dependencies, and a decoder to generate counterfactual sequences. As visualized in Figure 1, $\{X,Y\}\subseteq V$ represents the sets of all windows of time.
@@
We build on the Time Causal VAE (acciaio2024time), which provides both the architectural foundation and the theoretical justification to enforce causality in VAEs through recurrent networks such as Long-Short Term Memory (LSTM) (hochreiter1997long) and the Gated Recurrent Unit (GRU) (cho2014properties). Their framework relies on the Causal Wasserstein distance (cheridito2025optimal), implemented via bicausal couplings, to ensure that the latent dynamics respect the causal structure. In our work, we adopt this formulation but extend it by incorporating an explicit Directed Acyclic Graph (DAG) representation into the model (Figure 2), enabling the evaluation of counterfactual queries in addition to causal generative modeling.
@@ 2.1. Training
The encoder network $Q_{\phi}(U\mid V)$ maps the input time series to a latent space via a hierarchical structure. An initial feedforward network extracts preliminary features that are then processed by GRU layers to capture temporal dependencies. The encoder outputs the mean $\mu$ and log-variance $\log\sigma^{2}$ parameters of the latent distribution. Following the VAE framework (kingma2013auto), we apply the reparameterization trick to sample latent variables,
@@
As illustrated by the causal structure in Figure 2, $X_{t-1}$ has a direct effect on $Y_{t}$. When sampling from the inferred probability distribution $P_{\theta}(X\mid u)$, we employ the RealNVP (dinh2017density) transformation to model complex conditional distributions. Following our construction, a separate latent space is encoded for each marginal stochastic process (e.g., $X$ and $Y$). In the decoder, the output must be concatenated with the latent state at the previous time step, $U_{t-1}$, which allows flexible modification when performing interventions such as $P(Y\mid\mathrm{do}(x))$.
@@ 2.2. Counterfactual Generation
During counterfactual generation (Figure 3), we follow a three-step process (pearl2009causality):
@@ 3. Experiments
For the experiments, we used two autoregressive AR models, inspired by the Ornstein–Uhlenbeck process (doob1942brownian). This formulation was chosen due to its well-known stationary properties, its mean-reverting behavior, and its common use in modeling financial time series (bjork2009arbitrage). In addition, working in a controlled setting grants us access to the ground truth. We generate synthetic data for
@@ A.2. Counterfactual Modelling
The following condition is sufficient to guarantee counterfactual consistent estimations (pan2024counterfactual).

### v4

@@ 1. Introduction
Market generators (Kondratyev and Schwarz, 2019) are numerical techniques that rely on generative models for the purpose of synthetic market data generation. They leverage architectures such as generative adversarial networks (GANs) (Wiese et al., 2020; Li et al., 2020). For example, in market making, probabilistic forecasts can enhance trading strategies, allowing more profitable and risk-aware trading (Lalor and Swishchuk, 2025). Important attributes of synthetic financial time series are the preservation of stylized facts (Cont, 2001) and rough path signatures (Muca Cirone et al., 2024).
Historically, understanding of causality in times series was limited to Granger causality (Kleinberg and Mishra, 2009). Granger causality assumes that the frequency of data measurement matches the true causal frequency of the underlying physical process (Gong et al., 2017). However, making investment decisions requires a causal attribution of the premia to risk factors (de Prado, 2025). As such, research begins to incorporate causal reasoning and explainability into time series foundation models (e.g., through attention mechanisms) (Marconi, 2025; Robertson et al., 2025). For example, (Sokolov et al., 2025) demonstrated the utility of specifying and analyzing detailed causal models for financial markets for investment management. (Oliveira et al., 2024) investigated superior financial time series forecasting that leverages causality-inspired models to balance the trade-off between invariance to distributional changes and minimization of prediction errors.
Research in generative models went beyond vector auto regression (VAR) (Stock and Watson, 2001) to account for structural causality through the framework of structural causal models (SCMs) (Pearl, 2009). Examples of time series causal methods are SCIGAN (Bica et al., 2020b), Time Series Deconfounder (Bica et al., 2020a), and Causal Transformer (Melnychuk et al., 2022). Deep structural causal models (DSCMs) (Poinsot et al., 2024) enable a specific type of conditioned generation, counterfactuals, which are possible realistic scenarios that have not yet happened (Horvath et al., 2025). They are a promising method for promoting financial stress tests (Gao et al., 2018), risk management, scenario analysis, and backtesting (Harvey and Liu, 2015).
@@ 2. Methodology
We propose a causal market simulator building on top of previous research in neural causal models (NCMs) (Xia et al., 2023) and variational autoencoder (VAE) (Kingma, 2013) based causal representation learning (Schölkopf et al., 2021) in time series. Our neural causal model for time series (TNCM-VAE) [^1] https://github.com/thummd/tncm consists of three main components: an encoder to infer latent representations, a causal mapping module to handle dependencies, and a decoder to generate counterfactual sequences. As visualized in Figure 1, $\{X,Y\}\subseteq V$ represents the sets of all windows of time.
@@
We build on the Time Causal VAE (Acciaio et al., 2024), which provides both the architectural foundation and the theoretical justification to enforce causality in VAEs through recurrent networks such as Long-Short Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and the Gated Recurrent Unit (GRU) (Cho et al., 2014). Their framework relies on the Causal Wasserstein distance (Cheridito and Eckstein, 2025), implemented via bicausal couplings, to ensure that the latent dynamics respect the causal structure. In our work, we adopt this formulation but extend it by incorporating an explicit Directed Acyclic Graph (DAG) representation into the model (Figure 2), enabling the evaluation of counterfactual queries in addition to causal generative modeling.
@@ 2.1. Training
The encoder network $Q_{\phi}(U\mid V)$ maps the input time series to a latent space via a hierarchical structure. An initial feedforward network extracts preliminary features that are then processed by GRU layers to capture temporal dependencies. The encoder outputs the mean $\mu$ and log-variance $\log\sigma^{2}$ parameters of the latent distribution. Following the VAE framework (Kingma, 2013), we apply the reparameterization trick to sample latent variables,
@@
As illustrated by the causal structure in Figure 2, $X_{t-1}$ has a direct effect on $Y_{t}$. When sampling from the inferred probability distribution $P_{\theta}(X\mid u)$, we employ the RealNVP (Dinh et al., 2017) transformation to model complex conditional distributions. Following our construction, a separate latent space is encoded for each marginal stochastic process (e.g., $X$ and $Y$). In the decoder, the output must be concatenated with the latent state at the previous time step, $U_{t-1}$, which allows flexible modification when performing interventions such as $P(Y\mid\mathrm{do}(x))$.
@@ 2.2. Counterfactual Generation
During counterfactual generation (Figure 3), we follow a three-step process (Pearl, 2009):
@@ 3. Experiments
For the experiments, we used two autoregressive AR models, inspired by the Ornstein–Uhlenbeck process (Doob, 1942). This formulation was chosen due to its well-known stationary properties, its mean-reverting behavior, and its common use in modeling financial time series (Björk, 2009). In addition, working in a controlled setting grants us access to the ground truth. We generate synthetic data for
@@ A.2. Counterfactual Modelling
The following condition is sufficient to guarantee counterfactual consistent estimations (Pan and Bareinboim, 2024).